How many whole numbers between 99 and 999 contain the digit '1'?
Answer by ikleyn(52788)

.
One hundred numbers from 100 to 199  having '1' in the hundreds position


plus  8*10 = 80  other numbers from 210 to 919 of the table 


    210    310   .   910
    211    311   .   911
    212    312   .   912
    ...    ...   .   ...
    219    319   .   919

having '1' in the tens position,


plus 8*9 = 72  other numbers from 201 to 991 of the table 


    201    301   .   901
    221    321   .   921
    231    331   .   931
    ...    ...   .   ...
    291    391   .   991

having '1' in the units position and not having '1' in the tens position.


In all, there are  100 + 80 + 72 = 252 such numbers.